OpenStudy (anonymous):

PLEASE HELP!! solve the equation (e^x-81e^-x)/2 =9 which is the solution in terms of natural logarithms? x=__________ OR there is no solution what is the decimal approximation for the solution? x=______________ OR There is no solution

OpenStudy (anonymous):

where is the eqation?

OpenStudy (anonymous):

the equation is on the top

OpenStudy (anonymous):

I've simplified this equation to e^2x-18e^x-81=0 You understood how to?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

e^x=t t^2-18t-81=0 you are aware with ax^2+bx+c=0 and how to find it's roots. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well you will get two values of 2, which will be values of e^x. Then take the log on both sides, you'll get x.

OpenStudy (anonymous):

so do I use the quadratic formal to solve the equation?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Quadratic for the value of e^x. and then logarithm to obtain x.

OpenStudy (anonymous):

I got \[\frac{ 18\pm \sqrt{648} }{ -36 }\]

OpenStudy (anonymous):

is that right? how do you logarithm it?

OpenStudy (anonymous):

@NeetziD